### v1

# HA-VLN: A Benchmark for Human-Aware Navigation in Discrete–Continuous Environments with Dynamic Multi-Human Interactions, Real-World Validation, and an Open Leaderboard

Abstract
Vision-and-Language Navigation (VLN) systems often focus on either discrete (panoramic) or continuous (free-motion) paradigms alone, overlooking the complexities of human-populated, dynamic environments. We introduce a unified Human-Aware VLN (HA-VLN) benchmark that merges these paradigms under explicit social-awareness constraints. Our contributions include: (1) a standardized task definition that balances discrete–continuous navigation with personal-space requirements; (2) an enhanced human motion dataset (HAPS 2.0) and upgraded simulators capturing realistic multi-human interactions, outdoor contexts, and refined motion–language alignment; (3) extensive benchmarking on 16,844 human-centric instructions, revealing how multi-human dynamics and partial observability pose substantial challenges for leading VLN agents; (4) real-world robot tests validating sim-to-real transfer in crowded indoor spaces; and (5) a public leaderboard supporting transparent comparisons across discrete and continuous tasks. Empirical results show improved navigation success and fewer collisions when social context is integrated, underscoring the need for human-centric design. By releasing all datasets, simulators, agent code, and evaluation tools, we aim to advance safer, more capable, and socially responsible VLN research.***Project: https://ha-vln-project.vercel.app/
Figure 1: HA-VLN Navigation Scenario. An agent navigates environments with dynamic human activities, adjusting its path based on natural language instructions and real-time sensor observations to avoid collisions. Key positions (e.g., , ) align with instructional cues referring to specific human behaviors. When the agent encounters a bystander on the phone (, Decision A), it intelligently turns right to avert a potential collision. On the right, RGB and Depth observations illustrate the agent’s panoramic view preceding decisions A, B, and C, capturing its dynamic responses to nearby humans. [Best viewed in color and high resolution.]
11footnotetext: Equal contribution. Work done during internship at UW.22footnotetext: Corresponding author.
1 Introduction
Vision-and-Language Navigation (VLN) [3, 56] enables robots to understand multimodal instructions and navigate real or simulated spaces [15, 49]. While recent VLN systems have shown encouraging results in constrained domains, most focus exclusively on either discrete (panoramic) or continuous (free-motion) navigation, omitting the rich human dynamics and partial observability of realistic environments [4, 23, 54]. These omissions compromise real-world applicability, as embodied agents must handle unpredictable human motion, avoid collisions, and process instructions referencing socially salient cues.
Motivation & Gaps. Despite notable progress, three major challenges persist in today’s VLN research. First, social awareness remains underexplored: human participants in the scene are commonly overlooked or reduced to inert obstacles, preventing the agent from respecting personal space or reacting to bystanders’ activities (see Figure 1). Second, instruction complexity is not well captured in existing corpora. Commands such as “Go around the chair” rarely reflect real-world contexts like “Go around the person talking on the phone” [39]. Third, static-environment assumptions dominate, neglecting real-time re-planning when people traverse corridors or gather spontaneously. In practice, social navigation demands partial observability and dynamic route adjustment.
Human-Aware VLN. To tackle these limitations, Human-Aware VLN (HA-VLN) [31] explicitly integrates dynamic human interactions, driving agents to navigate among people engaged in various everyday behaviors. Although HA-VLN moves closer to realistic settings, it remains constrained by four main issues (Appendix Table A1): (1) Limited Human Motion Data, as initial datasets (e.g., HAPS 1.0 [31]) sometimes misalign textual labels and 3D motion sequences; (2) Discrete Navigation Bias, since many simulators restrict navigation to panoramic hops [28]; (3) Underdeveloped Multi-Human Interactions, typically modeling only a single person in indoor scenes; and (4) Instruction Generation Deficits, where automated text expansions omit time-varying human behaviors, limiting agents’ adaptability in dynamic, crowded environments.
Our Approach & Contributions. In this paper, we substantially extend HA-VLN by unifying discrete and continuous navigation, coupling high-fidelity human motion data, multi-agent interactions, and real-time rendering under a single umbrella. Concretely, we make four primary contributions, each elaborated in its respective section:
  • Standardized Task & Metrics [Sec. 2]. We unify discrete and continuous navigation under social-awareness constraints, ensuring consistent goals and evaluations.
  • Upgraded Data & HA-VLN Simulator [Sec. 3]. Building on HAPS 1.0 [31], we present HAPS 2.0 (486 SMPL sequences) and two advanced simulators (HA-VLN-DE, HA-VLN-CE) that incorporate multi-view human annotation, dual-thread rendering, and rigorous collision checks for up to 910 active individuals (Figure 2).
  • Comprehensive Benchmarking [Sec. 4]. We enrich R2R-CE with 16,844 human-centric instructions (Table A3) and benchmark multiple agents under unified metrics, unveiling challenges arising from multi-human dynamics and partial observability.
  • Real-World Validation & Leaderboard [Sec. 5]. We robustly demonstrate sim-to-real transfer using a physical robot successfully navigating crowded indoor areas, and provide a public leaderboard for comprehensive discrete–continuous evaluations in multi-human scenarios.
As visualized in Figures 1 and 2, introducing realistic human motion significantly amplifies task complexity, leading to pronounced performance drops in baseline VLN models. By bridging discrete and continuous navigation, enriching human-motion fidelity, and validating results both in simulation and on physical robots, our approach narrows the gap between conventional VLN setups and real-world, socially aware navigation systems. We release all related datasets, simulators, agent code, and evaluation tools to further fuel progress in this emerging research area.
Figure 2: HA-VLN-CE Simulator. We extend photorealistic Habitat environments by integrating diverse and dynamic human activities using the HAPS 2.0 dataset (172 activities, 486 motion models, 58,320 frames). The annotation pipeline proceeds in two stages: (i) a coarse-to-fine optimization approach with PSO and synchronized multi-view cameras, and (ii) a human-in-the-loop iterative refinement process to enrich multi-human interactions and movements. Real-time rendering uses a signaling mechanism to update human motion dynamically, enabling continuous collision detection and agent-environment interplay. [Best viewed in color and high resolution.]
2 Human-Aware VLN Task
Motivation & Overview. We present Human-Aware Vision-and-Language Navigation (HA-VLN), a next-generation VLN framework that specifically requires agents to navigate among dynamically moving people while following natural-language instructions. Unlike standard VLN—where humans are either absent or treated as static obstacles—HA-VLN compels agents to anticipate human motion, respect personal space, and adapt paths to avoid collisions. As shown in Figure 1, the agent must interpret instructions referencing ongoing activities (e.g., “Go upstairs where someone is pacing on the phone…”) and then maneuver accordingly in a photorealistic 3D space.
State & Action Spaces. At each time step $t$, the agent’s state is
$s_{t}\;=\;\langle\mathbf{p}_{t},\,o_{t},\,\Theta_{t}^{\text{FOV}}\rangle,$
where $\mathbf{p}_{t}$ is its 3D location, $o_{t}$ its orientation, and $\Theta_{t}^{\text{FOV}}$ its egocentric view. In discrete environments (DE), agents hop among pre-defined viewpoints, each providing an RGB observation. In continuous environments (CE), agents see an RGB+D feed with a $90^{\circ}$ field of view and can move in small increments (e.g., $0.25\,\mathrm{m}$ forward, $15^{\circ}$ turn). Across both settings, the action space is
$\mathcal{A}\;=\;\{\,a_{\text{forward}},\,a_{\text{left}},\,a_{\text{right}},\,a_{\text{up}},\,a_{\text{down}},\,a_{\text{stop}}\}.$
Figure 2 provides examples of how these actions evolve within complex, human-occupied scenes.
Human-Aware Constraints. Beyond reaching a goal, agents must obey strict social protocols:
  • Dynamic Human Models. People move naturally according to 3D motion trajectories from HAPS 2.0, which are continuously updated in real time.
  • Personal Space. Agents must avoid getting too close (under 3 m in DE or within the sum of radii in CE).
  • Human-Focused Instructions. Language often describes people and their activities (e.g., “Go around the person talking on the phone”), necessitating alignment between textual prompts and live visual inputs.
These detailed human placements and movements are annotated via a multi-stage pipeline (Sec. 3), entailing extensive labeling and verification to ensure realism.
Dynamic & Partial Observations. Because people can move unpredictably, each time step is a partially observable Markov decision process (POMDP). The new state $s_{t+1}$ depends on both the agent’s action and concurrent human movements (e.g., someone stepping aside or emerging from a corridor). Agents must infer unobserved factors—such as whether a person will yield space—and balance exploration (seeking alternative routes) and exploitation (continuing on a known path) to reach the target efficiently.
Challenges & DE–CE Synergies. Building socially compliant robots in crowded settings poses three key Sim2Real challenges: (i) Social Navigation, i.e., collision-free motion that respects personal space; (ii) Human-Aligned Instructions, where language may refer to fleeting activities or interactions; and (iii) Adaptive Re-Planning, as humans can block or unblock hallways unpredictably. While DE allows rapid prototyping via discrete viewpoint jumps, CE approximates real-world dynamics and motion fidelity. By integrating both approaches, HA-VLN spans the spectrum from large-scale simulation to practical on-robot deployment, significantly broadening research in socially aware, human-centric navigation.
3 HA-VLN Simulator
Motivation & Overview. The HA-VLN Simulator addresses long-standing challenges in socially-aware navigation by placing multiple, dynamically moving humans into both discrete and continuous 3D environments. While many existing simulators either overlook human behaviors or model them as static obstacles, our simulator features high-fidelity motion, multi-person interactions, and real-world complexities such as group gatherings, spontaneous movements, and personal-space constraints. Table A2 situates our approach relative to Habitat 3.0 [42] and HabiCrowd [48], highlighting broader scenario coverage, more natural motion types, and explicit instruction alignment. Meanwhile, Table A3 contrasts HAPS 2.0 with its predecessor, HAPS 1.0 [31], underscoring notable advancements in motion diversity, alignment accuracy, and overall realism. By building on HAPS 2.0, we leverage 486 motion sequences that span indoor and outdoor activities. We provide two complementary modules: HA-VLN-CE for continuous navigation and HA-VLN-DE for discrete navigation. Both modules share a unified API (Sec. 3), offering consistent access to human states, dynamic scene updates, and collision checks. Figure 2 presents an overview of how these elements converge within an agent’s action-observation loop.
HAPS 2.0 Dataset. Evolving from HAPS 1.0 [31], HAPS 2.0 encompasses 486 3D human-activity descriptions covering 26 region types (e.g., living rooms, gyms, sidewalks). Each description (e.g., “workout gym exercise: An individual running on a treadmill”) is validated through human surveys and ChatGPT-4 [6] for plausibility. We employ the Motion Diffusion Model (MDM) [17] to transform these text descriptions into 120-frame SMPL mesh sequences, $\mathcal{H}=\langle h_{1},\ldots,h_{120}\rangle$, capturing detailed 3D pose and shape information. Figure A2 illustrates various indoor and outdoor contexts, while Figure A3 depicts representative motions (e.g., climbing stairs, running).
Annotation Pipeline: Coarse-to-Fine. To position these motions in large-scale 3D scans, we employ a two-stage annotation procedure. Stage 1 (Coarse) uses Particle Swarm Optimization (PSO) [25] (Algorithm A1) to allocate motions within each region $\mathbf{R}$, bounded by $\mathbf{B}_{\text{lo}},\mathbf{B}_{\text{hi}}$, and link them to suitable objects $\mathbf{O}$. A subset $\mathbf{H}^{\prime}\subseteq\mathbf{H}$ is filtered by region labels, then paired with objects via semantic cues (e.g., “sitting on a chair” aligns with a chair object). PSO penalizes boundary violations and personal-space intrusions (1 m), producing highly plausible initial placements. Stage 2 (Fine) refines these placements with a precise nine-camera approach [22, 40] (Figure A1). Annotators iteratively adjust rotations and positions to eliminate collisions or interpenetrations (e.g., limbs intersecting walls), requiring over 430 hours of human labeling to achieve 529 precisely aligned models across 90 scans. To further enhance diversity, we integrate a human-in-the-loop process [11] for generating multi-person scenarios: large language models propose contextual interactions (Listings LABEL:list:dataset_prompt and LABEL:list:few-shot-example), which annotators then refine. Figure A3 illustrates enhanced interactions (e.g., group discussions). After both stages, we assemble 910 human models distributed over 428 regions in 90 scans, including 268 complex motions (e.g., running, stair climbing). Figure A5 and Appendix B.8 detail their distributions, revealing extensive coverage of diverse motion types and scene layouts.
Figure 3: Motion Analysis. (a) Top three motions from Stage 1 (without enrichment) and Stage 2 (with enrichment). (b) Overall statistics of human activities, comparing interaction types, movement distances, and the number of models used. Enrichment extends both the variety and dynamic aspects of human activities.
Real-Time Rendering & Agent Interaction. Our simulator updates human motions in real time rather than relying on static captures. A dual-thread producer-consumer architecture (Algorithm A2) orchestrates frame updates, with Thread 1 enqueuing “refresh” signals and Thread 2 processing them in sync with the agent’s action cycle. Each motion spans up to 120 frames; once Thread 2 detects a signal, it discards old meshes and loads new ones, maintaining retrieval delays under 50 ms. Figure A2 exemplifies how multiple humans can simultaneously populate a shared environment. Agents perceive these moving entities through a navigation mesh (navmesh) [45]. Collisions are flagged if the distance between bounding volumes is less than their combined radii, leading to an automatic revert. This mechanism ensures agents learn to respect personal space and navigate effectively around dynamic crowds.
Discrete vs. Continuous Settings. HA-VLN-CE (Continuous) allows agents to move in real-valued increments (e.g., $0.25\,\mathrm{m}$ forward, $15^{\circ}$ turns), supporting fine-grained collision avoidance and social reactivity. As Figure A5 shows, each scene can contain up to 10 humans, with simulation speeds reaching 30–60 FPS on standard GPUs. HA-VLN-DE (Discrete) extends HA3D [31] to incorporate HAPS 2.0 data, encompassing both indoor and outdoor settings. Here, agents hop among panoramic viewpoints while humans continue to move, preserving core social-navigation challenges. We map continuous positions to discrete viewpoints [31], add small offsets to refine alignment, and integrate 627 annotated humans across 90 buildings.
Unified API.  We provide a unified API for both modes with three core functionalities: (1) Human State Queries retrieve bounding volumes, motion frames, and semantic annotations of nearby humans; (2) Dynamic Scene Updates notify agents of newly moved humans or environmental changes; and (3) Collision Checks detect whether a proposed move (e.g., forward step or viewpoint hop) would result in intersection with a human. By unifying HAPS 2.0 with coarse-to-fine annotation, real-time multi-human rendering, and a single API across discrete and continuous settings, our HA-VLN Simulator significantly broadens the scope of socially-aware navigation research. Figures A2, A3, and A5 depict the simulator’s capacity to capture rich human behaviors, while Tables A1 and A3 underscore our advantages over earlier simulators, including Habitat 3.0 and HabiCrowd, as well as the improvements of HAPS 2.0 relative to HAPS 1.0. Appendix B.7 outlines environment sizes, latency considerations, and usage examples.
Figure 4: HA-R2R Dataset Analysis. (a) Instruction-length distribution by human group size (no humans to more than three). (b) Comparison of instruction lengths between HA-R2R and R2R-CE.
4 HA-VLN Agents
This section introduces our Human-Aware Room-to-Room (HA-R2R) dataset and two navigation agents: HA-VLN-VL and HA-VLN-CMA. We aim to endow agents with enhanced understanding of social cues—respecting personal space, avoiding collisions, and reacting to bystanders. As illustrated in Figures A8 and A9, dynamic human behaviors impose additional complexity and variability, necessitating robust perception and policy strategies.
HA-R2R Dataset. Although the Room-to-Room in Continuous Environment (R2R-CE) dataset [28] supports continuous navigation, it lacks explicit focus on human interactions. We therefore develop HA-R2R, which extends R2R-CE with 16,844 carefully curated instructions highlighting social nuances (e.g., conversations, corridor crossings, near-collision events). Table A4 showcases representative directives, while Figure A6 reveals the expanded vocabulary centered around human-centric language.
To generate these enriched instructions, we design targeted prompts for large language models (Appendix C.2), capturing a diverse range of real-world social scenarios. This augmentation marks a significant leap from static paths to ones requiring agents to interpret, for example, “avoid the couple chatting near the bar” or “yield to someone crossing the hallway.” Our comparative analysis between baseline and augmented instructions (Appendix C.3) highlights both the extensive annotation workload and the potential of HA-R2R to advance human-aware navigation research.
HA-VLN-VL Agent. HA-VLN-VL agent focuses on visual-language alignment when encountering socially enriched instructions. Adapted from Recurrent VLN-BERT [20], it discards actor-critic methods (e.g., A2C [27]) in favor of a simpler imitation learning approach, thereby emphasizing how stronger multimodal grounding alone can elevate performance in complex settings.
At each timestep $t$, the agent updates its hidden state $s_{t}$ and predicts an action distribution $p_{t}^{a}$:
$s_{t},\;p_{t}^{a}\;=\;\text{HA-VLN-VL}(s_{t-1},\,X,\,V_{t}),$
where $X$ is the tokenized instruction (often referencing multiple humans or group interactions), and $V_{t}$ encodes the fused RGB-depth view. A multi-layer Transformer processes these inputs, using a specialized state token to attend to language and visual tokens. Final action probabilities result from an average of attention weights:
$p_{t}^{a}\;=\;\overline{\text{AveragePool}}_{s,v}^{l}.$
By fine-tuning from Prevalent [20] on HA-R2R, HA-VLN-VL concentrates on the semantic interplay between social directives and the observed scene. Further architectural details appear in Appendix C.6.
HA-VLN-CMA Agent. While HA-VLN-VL excels at interpreting nuanced instructions, HA-VLN-CMA prioritizes collision avoidance & real-time adaptation in human-filled environments. Drawing on cross-modal attention (CMA) [28], it fuses textual embeddings $l=\text{BERT}(I)$ and visual features $v_{t}=\text{ResNet}(o_{t})$ each timestep. A multi-head attention mechanism produces a combined representation $f_{t}$, which an MLP maps to:
$P(a_{t}\mid f_{t})\;=\;\text{Softmax}\!\bigl{(}\text{MLP}_{\mathrm{action}}(f_{t})\bigr{)}.$
Figure A7(b) outlines the pipeline, while Appendix C.7 discusses implementation details. Human-populated settings entail partial observability and unpredictable motion. To address these challenges, we leverage Environmental Dropout (Envdrop) [46] & Dataset Aggregation (DAgger) [44]. Envdrop randomly masks features in the visual-language streams, simulating occlusions by crowds or obstructing objects. DAgger maintains a replay buffer of agent states under both expert and learned policies, enabling the agent to iteratively correct prior errors. This additional training effort improves trajectory re-planning when confronted with obstacles or unexpected human movements.
Algorithm 1 HA-VLN-CMA Training Procedure
0:  Expert policy $\pi^{*}$, parameters $\theta$, replay buffer $B$
0:  Trained CMA model parameters $\theta$
1:  Initialize parameters $\theta$ (e.g., via imitation learning)
2:  for iteration $=1$ to $N$ do
3:     Execute policy $\pi_{\theta}$ to collect rollouts
4:     With probability $\beta$, follow expert $\pi^{*}$
5:     Store $(s_{t},a_{t},r_{t})$ in $B$
6:     Apply Envdrop to $(v_{t},l)$
7:     $\theta\leftarrow\theta-\eta\nabla_{\theta}\,\mathcal{L}_{\mathrm{DAgger}}(B,\theta)$
8:  end for
Analysis & Insights. Figures A8 & A9 demonstrate how our agents respond to bystanders, underscoring that collision likelihood and route deviation both significantly increase when humans crowd key passageways. HA-VLN-CMA re-plans aggressively if a person suddenly blocks its immediate path, whereas HA-VLN-VL draws upon textual grounding to maintain a polite distance from social gatherings. This distinction showcases our dual innovation: a socially enriched dataset (HA-R2R) and specialized models that jointly facilitate robust human-aware navigation. Table A4 & Figure A6 further illustrate the magnitude and complexity of our annotations, emphasizing the extensive workload in data preparation, model adaptation, and thorough validation. Finally, in Section 6, we present comprehensive evaluations on continuous & discrete tasks derived from HA-R2R, illustrating the complementary strengths of HA-VLN-VL (strong language-vision fidelity) and HA-VLN-CMA (robust collision avoidance).
Figure 5: Human-aware navigation with multiple bystanders. Top: The robot receives instructions emphasizing caution as one person moves from the lounge to the plant while another prepares to exit. Middle: A third-person view highlights the robot’s trajectory and path planning among dynamic bystanders. Bottom: The onboard camera shows the robot’s egocentric observations.
5 Real-World Validation & Leaderboard
Real-World Validation & Setup. We deploy our trained agents on a Unitree Go2-EDU quadruped, equipped with an Intel Realsense D435i RGB-D camera, MID360 3D LiDAR, and IMU for onboard perception and control. As Figure 5 illustrates, experiments are conducted in four indoor spaces (office, living room, hallway, lobby), each populated by 2–4 free-moving volunteers. The agent navigates safely in moderately congested conditions but faces challenges in tight corridors or sudden crowd convergence, highlighting the need for robust re-planning under partial observability.
HA-R2R Test Dataset & Leaderboard. Building on R2R-CE, we present HA-R2R, featuring 16,844 instructions across 90 building scans with 910 annotated human models (see Sections 34). The test partition—3,408 instructions in 18 withheld buildings—intentionally emphasizes multi-human routes. While retaining path continuity from R2R-CE, we introduce refined goals to stress social awareness. The dataset is split into training (10,819), seen validation (778), unseen validation (1,839), and testing (3,408). We also host leaderboards for HA-R2R-DE (discrete) and HA-R2R-CE (continuous), focusing on both collision-related (TCR, CR) and navigation (NE, SR) metrics. Submissions may include agent code or trajectories, providing reproducible, server-side evaluations and setting a new benchmark for human-centric, dynamic VLN research.
Table 1: HA-VLN-CE Results Across Validation (Seen/Unseen) and Test Splits. “HA-VLN-CMA” denotes the full version of HA-VLN-CMA (+DA +EV). Metrics include NE (Navigation Error, meters), TCR (Total Collision Rate), CR (Collision Rate per step), and SR (Success Rate), with lower NE/TCR/CR and higher SR indicating better performance.
Agent Validation Seen Validation Unseen Test
Retrained Zero-shot Retrained Zero-shot Retrained Zero-shot
[…]
HA-VLN-CMA-Base 7.63 63.09 0.77 0.05 7.88 63.84 0.75 0.04 7.34 47.06 0.77 0.07 7.95 63.96 0.76 0.03 7.30 47.55 0.76 0.07 7.89 62.14 0.74 0.04
HA-VLN-CMA-DA 6.11 17.45 0.61 0.17 6.95 37.85 0.72 0.07 7.00 27.25 0.69 0.09 7.05 38.22 0.73 0.05 7.12 28.33 0.69 0.08 6.98 36.53 0.73 0.06
HA-VLN-CMA 5.61 3.34 0.60 0.17 7.10 29.99 0.69 0.11 6.23 8.10 0.69 0.10 6.62 32.48 0.70 0.09 6.64 9.23 0.72 0.09 7.09 31.80 0.75 0.09
HA-VLN-VL 5.02 4.44 0.52 0.20 7.82 3.67 0.45 0.05 5.35 6.63 0.59 0.14 7.15 3.97 0.46 0.06 5.52 5.96 0.63 0.14 7.41 3.38 0.58 0.07
BEVBert [1] 5.53 3.64 0.46 0.27 6.11 4.29 0.47 0.19 5.51 4.71 0.55 0.21 6.10 5.72 0.56 0.15 6.33 4.25 0.58 0.18 6.54 4.39 0.54 0.14
ETPNav [2] 5.17 4.07 0.43 0.24 7.72 6.31 0.61 0.12 5.43 6.94 0.58 0.17 7.40 7.94 0.71 0.08 5.94 5.96 0.58 0.16 7.59 5.64 0.73 0.09
Table 2: Performance in Unseen Environments (HA-VLN-CE) with Relative Gaps. We highlight $\dagger$” for zero-shot models (trained on VLN-CE). Columns for NE/TCR/CR/SR are available in Table 1, so here we focus on relative improvements (%). Entries show changes vs. the VLN-CE baseline.
Agent Retrained Gap (%) Zero-Shot Gap (%)
[…]
HA-VLN-CMA-Base -9.8 -69.6
HA-VLN-CMA-DA -13.8 -66.7
HA-VLN-CMA -10.0 -68.8 +10.2 -71.9
HA-VLN-VL -6.8 -68.2 +24.6 -86.4
BEVBert [1] +20.6 -64.4 +33.5 -74.6
ETPNav [2] +15.3 -70.2 +57.1 -86.0
Table 3: Retrained Performance on HA-VLN-DE (Unseen Environments). NE: Navigation Error (meters, lower is better), SR: Success Rate (higher is better), TCR: Total Collision Rate (lower is better), CR: Collision Rate per Step (lower is better).
Agent VLN HA-VLN-DE
[…]
Speaker-Follower [12] 6.62 0.35 7.44 0.32 0.72 0.21
Rec (PREVALENT) [20] 3.93 0.63 6.12 0.29 0.81 0.33
Rec (OSCAR) [20] 4.29 0.59 6.12 0.28 0.78 0.34
Airbert [16] 4.01 0.62 5.54 0.30 0.83 0.36
Table 4: Impact of Human Presence (hp) and Interaction Enrichment (enrich). We evaluate both continuous (CE) and discrete (DE) settings, reporting NE: Navigation Error, TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate.
[…]
CE BEVBert 6.10 5.72 0.56 0.15
CE ETPNav 7.40 7.94 0.71 0.08
DE Rec (PREVALENT) 7.31 0.31 0.79 0.22
CE BEVBert 6.32 ($\uparrow$3.6%) 5.11 ($\downarrow$10.7%) 0.46 ($\downarrow$17.9%) 0.17 ($\uparrow$13.3%)
CE ETPNav 7.35 ($\downarrow$0.6%) 6.12 ($\downarrow$22.9%) 0.63 ($\downarrow$11.3%) 0.10 ($\uparrow$25.0%)
DE Rec (PREVALENT) 7.52 ($\uparrow$2.9%) 0.27 ($\downarrow$12.9%) 0.64 ($\downarrow$19.0%) 0.27 ($\uparrow$22.7%)
CE BEVBert 6.13 ($\uparrow$0.5%) 3.25 ($\downarrow$43.2%) 0.35 ($\downarrow$37.5%) 0.19 ($\uparrow$26.7%)
CE ETPNav 7.75 ($\uparrow$4.7%) 4.47 ($\downarrow$43.7%) 0.53 ($\downarrow$25.4%) 0.14 ($\uparrow$75.0%)
DE Rec (PREVALENT) 7.33 ($\uparrow$0.3%) 0.19 ($\downarrow$38.7%) 0.42 ($\downarrow$46.8%) 0.26 ($\uparrow$18.2%)
Table 5: Cross-Evaluation of BEVBert (CE) vs. Rec (PREVALENT) (DE). Each model is trained/validated under two different instruction sets (HA-R2R vs. R2R/R2R-CE). NE: Navigation Error (m, lower is better); SR: Success Rate (higher is better). The gray cells () indicate performance changes when models are trained on R2R/R2R-CE instructions but validated on HA-R2R.
Env Training Validation Val (Unseen)
[…]
CE HA-VLN-CE HA-R2R HA-VLN-CE HA-R2R 5.51 0.21
R2R-CE HA-R2R 6.23 ($\uparrow$13.1%) 0.15 ($\downarrow$28.6%)
DE HA-VLN-DE HA-R2R HA-VLN-DE HA-R2R 5.01 0.39
R2R HA-R2R 6.11 ($\uparrow$22.0%) 0.24 ($\downarrow$38.5%)
Table 6: Impact of Step Size on Navigation. Increasing the step size transitions from finer-grained (CE-like) control to more discrete steps. We show results for both BEVBert [1] and ETPNav [2] on seen/unseen validations. NE: Navigation Error (m), TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate. The default step size for CE is 0.25 m.
Agent Step Size Validation (Seen) Validation (Unseen)
[…]
BEVBert[1] 0.10 5.65 8.43 0.50 0.23 5.41 12.60 0.54 0.22
0.25 (CE Default) 5.53 3.64 0.46 0.27 5.51 4.71 0.55 0.21
0.40 5.60 1.77 0.39 0.28 5.63 2.63 0.44 0.25
1.00 5.82 0.42 0.21 0.29 5.54 0.63 0.26 0.26
2.25 7.66 0.09 0.10 0.03 7.23 0.10 0.10 0.03
ETPNav[2] 0.10 5.15 11.70 0.54 0.20 5.47 18.66 0.64 0.16
0.25 5.17 4.07 0.43 0.24 5.43 6.94 0.58 0.17
0.40 5.11 2.43 0.36 0.26 5.32 3.77 0.46 0.21
1.00 6.67 0.49 0.25 0.24 6.76 0.79 0.32 0.17
2.25 7.61 0.10 0.10 0.02 7.21 0.13 0.12 0.03
Table 7: Ablation on RGB/Depth Inputs. We compare BEVBert [1] and ETPNav [2] on seen/unseen validations. denotes the sensor is enabled, while is disabled. Gray cells highlight relative performance changes (in %) upon removing or adding a modality. NE: Navigation Error (m), TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate. Best viewed in color.
Agent RGB Depth Validation (Seen) Validation (Unseen)
[…]
BEVBert[1] 6.23 ($\uparrow$12.6%) 4.55 ($\uparrow$25.0%) 0.49 ($\uparrow$6.5%) 0.19 ($\uparrow$29.6%) 5.79 ($\uparrow$5.1%) 4.97 ($\uparrow$5.5%) 0.53 ($\uparrow$3.6%) 0.15 ($\downarrow$28.6%)
5.68 ($\uparrow$2.7%) 3.77 ($\uparrow$3.6%) 0.48 ($\uparrow$4.3%) 0.25 ($\downarrow$7.4%) 5.50 ($\uparrow$0.2%) 4.73 ($\uparrow$0.4%) 0.53 ($\uparrow$3.6%) 0.20 ($\downarrow$4.8%)
5.53 3.64 0.46 0.27 5.51 4.71 0.55 0.21
ETPNav[2] 6.14 ($\uparrow$18.8%) 6.07 ($\uparrow$49.1%) 0.56 ($\uparrow$30.2%) 0.17 ($\uparrow$29.2%) 6.38 ($\uparrow$17.5%) 7.44 ($\uparrow$7.2%) 0.65 ($\uparrow$12.1%) 0.13 ($\downarrow$23.5%)
4.92 ($\uparrow$4.8%) 5.45 ($\uparrow$33.9%) 0.55 ($\uparrow$27.9%) 0.21 ($\downarrow$12.5%) 5.94 ($\uparrow$9.4%) 7.23 ($\uparrow$4.2%) 0.65 ($\uparrow$12.1%) 0.16 ($\downarrow$5.9%)
5.17 4.07 0.43 0.24 5.43 6.94 0.58 0.17
Table 8: Step Size Combination Performance Across Scenarios. NE: Navigation Error (m), TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate.
[…]
1.00 6.85 26.97 0.94 0.004
2.25 8.79 112.78 0.97 0.000
Table 9: Navigation Success Rate (NSR, %) in four environments (Living Room, Office, Hallway, and Lobby) with or without human presence (w/o vs. w/), averaged over 10 trials. Bold numbers indicate the highest performance in each column.
Method Living Rm Office Hallway Lobby Overall
w/o w/ w/o w/ w/o w/ w/o w/ w/o w/
HA-VLN-CMA-Base 0.2 0.1 0.2 0.1 0.2 0.1 0.2 0.1 0.2 0.1
HA-VLN-VL 0.4 0.1 0.4 0.1 0.5 0.2 0.4 0.1 0.4 0.1
6 Experiments
We evaluate our methods in two primary settings: (i) human-populated simulations within HA-VLN-CE (continuous) and HA-VLN-DE (discrete), and (ii) real-world robot deployments. All evaluations use four metrics—Navigation Error (NE), Success Rate (SR), Total Collision Rate (TCR), and Collision Rate (CR)—to measure navigation accuracy and social compliance. Extended formulas appear in Appendix D.1.
6.1 HA-VLN-CE: Continuous Navigation
Task Definition. HA-VLN-CE extends VLN-CE [28] by populating photorealistic 3D environments with multiple, independently moving humans. Agents navigate continuously (forward, turn) based on natural-language instructions, aiming to reduce NE while limiting collisions (TCR, CR). This setup mirrors real-world conditions where bystanders can alter their paths unexpectedly, demanding reactive policies and sophisticated sensory integration.
Baselines. We systematically benchmark two notable continuous-navigation models, BEVBert [1] and ETPNav [2], together with our HA-VLN-CMA and HA-VLN-VL agents. Each approach is evaluated under two configurations: Retrained, which leverages HA-R2R’s multi-human data, and Zero-shot, which relies solely on R2R-CE training. Contrasting these configurations further underscores the importance of explicit, human-centric instruction for truly robust social navigation.
Results. Table 1 shows pronounced gains when models incorporate HA-R2R data. For instance, BEVBert’s TCR drops from 4.29 to 3.64 in seen environments and from 5.72 to 4.71 in unseen ones. In stark contrast, Zero-shot models frequently collapse in high-density scenarios, with TCR surpassing 30 if they have no exposure to human-centric interactions. Table 2 further demonstrates that omitting multi-human data can diminish SR by up to 70% in unseen splits, underscoring the critical role of socially aware training for both collision avoidance and trajectory completion. These improvements suggest that explicit references to bystanders and dynamic crowd behavior substantially enhance real-world navigational readiness.
Ablation Studies. (1) Discrete vs. Continuous. Table 3 finds that discrete agents (HA-VLN-DE) can achieve moderate SR yet suffer high collisions in crowded scenes, implying simpler viewpoint-based control is inadequate for dynamic social contexts. (2) Human Presence. Table 4 shows how higher bystander density escalates TCR, underscoring the need for more robust collision avoidance in large crowds. (3) Instruction Alignment. Table 5 reveals that replacing HA-R2R with standard R2R-CE text raises NE by 13–22% and cuts SR by up to 38%, confirming the value of explicit human-motion references. (4) Step Size. Tables 6 and 8 show finer increments refine crowd responses but may inflate TCR via excess micro-adjustments. (5) Sensor Modalities. Table 7 confirms that adding depth consistently lowers collisions and raises SR, reflecting the importance of 3D cues for navigating around moving bystanders.
Insights. Together, these findings confirm that continuous, human-aware navigation benefits significantly from training on multi-human data, realistic instructions referencing bystander behavior, and richer sensor inputs. Failure to address any of these factors can lead to more collisions and lower success, highlighting the multi-faceted challenges of real-world social navigation.
6.2 HA-VLN-DE: Discrete Navigation.
Task Definition. HA-VLN-DE extends panoramic VLN [3] by placing multiple moving individuals in a viewpoint-based environment. Although discrete control restricts possible movements to predefined locations, it does not eliminate collisions, as bystanders may occupy a chosen viewpoint or block key transitions in narrow corridors.
Results. Table 3 compares top discrete agents on multi-human tasks. While Airbert achieves a moderate SR=0.36, it can incur a CR up to 0.83, illustrating persistent collision risks. Approaches that overlook human dynamics or personal space often fail when multiple bystanders converge, particularly in tight junctions or doorways. Consequently, viewpoint-based navigation alone is insufficient for robust social navigation—adaptive collision-avoidance strategies remain essential even in discrete settings.
6.3 Real-World Experiments.
Setup. We evaluate our agents in four everyday indoor environments—office, living room, hallway, and lobby—under two conditions: (i) no bystanders and (ii) 2–4 free-moving volunteers. This setup simulates realistic indoor traffic patterns and partial observability.
Observations. As illustrated in Figure 5, the robot frequently pauses or yields to avoid oncoming pedestrians. In the absence of bystanders, it navigates smoothly (Figure A11), but collisions arise in cramped corridors or when crowds converge suddenly (Figure A12). We observe similar patterns in living-room environments (Figures A13A14) and hallways (Figure A15).
Table 9 shows the average NSR (Navigation Success Rate) across 10 trials in each environment. While human presence invariably lowers NSR, HA-VLN-VL consistently outperforms HA-VLN-CMA-Base, demonstrating stronger adaptability to dynamic motion. Still, partial observability and abrupt group formations remain challenging, especially in narrow passages or at congested junctions. Additional metrics—such as collision frequencies and path divergence—appear in Appendix D.4, where we further detail performance under varying crowd densities.
Overall, these experiments confirm that simulation-trained, multi-human navigation policies can indeed transfer to physical robots. However, further refinement in collision forecasting and reactive control is needed to handle unpredictable human behavior in tight indoor settings.
7 Conclusion
We propose a Human-Aware VLN (HA-VLN) framework unifying discrete and continuous navigation under realistic multi-human conditions. By integrating dynamic human motion, refined annotations, and high-fidelity simulators, our HA-R2R dataset emphasizes human-centric instructions. Experiments show social awareness, multi-human interactions, and partial observability greatly increase complexity, reducing advanced agents’ performance. Nevertheless, our approach balances safety, efficiency, and personal space. Real-world tests confirm sim-to-real transfer, while our public leaderboard standardizes evaluations. By releasing all data, simulators, agents, and tools, we promote socially responsible, context-aware navigation in dynamic, human-populated environments.
Appendix
This supplementary material provides expanded details and results that complement the main paper. Section A offers a comprehensive literature survey focusing on three key research challenges. Section B describes our dataset construction, annotation protocols, real-time rendering methods, API design, and additional insights on annotation data. Section C presents an in-depth overview of the HA-R2R dataset and the proposed navigation agents. Finally, Section D includes detailed evaluation metrics, additional numerical results, visualized navigation outcomes, and real-world robot validation studies, each supplemented with thorough analysis. For further resources, the project page is accessible at https://ha-vln-project.vercel.app/, and the datasets are provided at DATASETS.
Appendix A Related Work
This appendix surveys the evolution of Vision-and-Language Navigation (VLN) tasks, simulators, and agent designs, with particular attention to how Human-Aware VLN (HA-VLN) advances the state of the art. We focus on three key aspects deemed critical for bridging the Sim2Real gap: (1) Socially Compliant Navigation, (2) Human-Aligned Instructions and Visual Cues, and (3) Dynamic Environments with Human Activities and Interactions. Table A1 summarizes how prior work compares under these dimensions.
A.1 Development of VLN Tasks
Early VLN tasks focused on basic indoor navigation—exemplified by Room-to-Room (R2R) [3, 15, 29]—and outdoor tasks like TOUCHDOWN [7] and MARCO [37]. Later efforts such as REVERIE [43] and VNLA [38] introduced object-centric or goal-driven navigation. While these approaches expanded the range of tasks, they typically overlooked real human behavior and social contexts. Dialogue-based tasks (e.g., DialFRED [13], CVDN [47]) incorporated interactive elements but did not account for dynamically moving bystanders or social-distance constraints. Initiatives like VLN-CE [28] moved closer to real-world conditions by enabling continuous (free-motion) navigation, yet remained largely devoid of explicit human factors [21, 29, 38, 47]. HA3D [31] addressed human motion and included human-oriented instructions but did not require agents to conform to social norms—e.g., maintaining safe distances or refraining from disturbing ongoing activities. Our proposed HA-VLN addresses these gaps by embedding all three essential elements—socially compliant navigation, human-referenced instructions, and dynamic human activities—into a single framework. Agents must plan routes among unpredictable bystanders, interpret language mentioning people and their behaviors, and uphold social standards. This integrated setup results in a benchmark that more closely aligns with real-world navigation demands.
A.2 Simulators for VLN Tasks
A reliable simulator is essential for developing and evaluating VLN agents. Early simulators like Matterport3D [3] and House3D [52] offered photorealistic or synthetic indoor environments but lacked mobile humans. Others, such as AI2-THOR [26] and Gibson [53], introduced more interactive elements yet typically assumed static or purely synthetic contexts, thus limiting their applicability for studying social compliance. Google Street View, used in some outdoor navigation tasks, presents static imagery with occasional humans in the scene but lacks dynamic or interactive elements. HA3D [31] moved a step further by including human activities and instructions referencing people, though it did not mandate socially compliant navigation. HabiCrowd [48] integrated crowds into photorealistic domains, improving visual diversity but omitting human-aligned instructions. Similarly, Habitat 3.0 [45] provides high-performance simulation without extensive multi-human or social-compliance features.
By contrast, our HA-VLN Simulator unifies dynamic human activities, photorealistic rendering, and social-compliance requirements. Agents perceive and react to evolving bystander behaviors—such as avoiding collisions or maintaining personal space—using both discrete (panoramic) and continuous (free-motion) navigation. As shown in Table A2, HA-VLN surpasses prior simulators in terms of scene count, motion diversity, and instruction alignment, providing a comprehensive platform for training and testing socially aware VLN agents. Specifically, we introduce 675 scenes (across 90 scenarios), 122 motion types, and a cohesive framework that supports instruction-driven dynamic human interactions. By supporting both discrete and continuous action spaces, HA-VLN further broadens its potential for addressing diverse navigation goals and real-world deployment challenges.
Table A2: Comparison with Habitat 3.0 and HabiCrowd. Our environment extends beyond existing simulators by offering a larger variety of natural motion types, an increased number of scenario types, and explicit instruction alignment for human-aware navigation. “Num Scenes” and “Scenario Types” refer to the total annotated scans and the distinct activities or region annotations, respectively; “Natural Motions” indicates whether human motions are photorealistically rendered and not purely script-based; “Scene & Interaction” denotes whether dynamic multi-human interactions are included in the environment.
Simulator Habitat Series Num Scenes Scenario Types Natural Motions Motion Type Scene & Interaction Instr. Align
Habitat 3.0 [41] 3 59 59 Yes 2 Yes No
HabiCrowd [48] 2 480 No None No No
Ours 1 675 90 Yes 122 Yes Yes
A.3 Agents for VLN Tasks
From early attention-based and reinforcement-learning approaches [36, 43, 50] to modern vision-language pre-training [35, 18, 32], VLN agents have grown increasingly adept at parsing instructions and navigating complex environments. However, most existing solutions—including EnvDrop [46], PREVALENT [18], and VLN-BERT [20]—rely on panoramic navigation, streamlining the action space but limiting realism of their movement. Recent efforts like NavGPT [57] and NaVid [55] explore continuous, egocentric navigation in partially dynamic worlds, yet they still lack explicit attention to human-aligned instructions or social compliance. In particular, these agents may not recognize the need to maintain safe distances, avoid disturbing activities, or adapt routes with active bystanders. Our HA-VLN agents address these gaps by navigating among multiple, moving humans and adhering to social norms. They interpret fine-grained, human-centric instructions and leverage visual cues that reflect real-world interactions, ensuring collision-free, respectful travel. This fusion of social compliance and dynamic human contexts sets HA-VLN apart, aligning agent behavior more closely with authentic, real-world challenges.
Table A3: Comparison of HAPS 1.0 vs. HAPS 2.0. We show the total number of motion categories, average accuracy and compatibility scores (both on a 1–10 scale), the number of failure cases (e.g., severe motion-description mismatches), and total annotation time. HAPS 2.0 features more diverse motions, improved motion-env alignment, and reduced failures, albeit at higher annotation effort.
Datasets Motions $\uparrow$ Accuracy (1-10) $\uparrow$ Compatibility (1-10) $\uparrow$ Failure Cases $\downarrow$ Annotation Time (hours)
HAPS 1.0 [31] 435 6.3 5.9 120 320 (verified by [31] authors)
HAPS 2.0 (ours) 486 8.5 8.1 0 430+
Appendix B Simulator Details
B.1 HAPS Dataset 2.0
We develop HAPS 2.0 to address the shortcomings of its predecessor, particularly in terms of mismatches between textual descriptions and motion data, as well as the limited diversity of region–motion associations.
Motion–Description Alignment. The original HAPS dataset contains 435 motion categories, each defined by a region (e.g., hallway) and a textual description (e.g., “Someone talking on the phone while pacing”). However, more than half of these pairs do not match accurately. We therefore conduct a two-round manual verification, where multiple volunteers determine whether each pair is valid. Motions that fail both rounds are removed, yielding 172 precisely aligned motions.
Diversifying Region–Motion Relationships. In the initial dataset, each region was tied to only a few rigidly defined motions (e.g., hallway mostly involves “pacing on a phone,” stairs focuses on “sliding down a banister” or “decorating the stairway”). Such narrow mappings cause biases and limit the realism of agent navigation. To remedy this, we reorganize region–motion associations, adapting the same motion to fit various environments, including both indoor and outdoor scenes. For instance, “talking on the phone” is re-contextualized to reflect whether someone is pacing upstairs or moving around a meeting room. This broader approach offers more faithful representations of human behavior and reduces environmental biases, thus improving real-world applicability.
HAPS 2.0 vs. HAPS 1.0. Table A3 quantitatively contrasts HAPS 2.0 with HAPS 1.0. We recruit 26 volunteers to evaluate every motion in both datasets on two 1–10 scales (motion accuracy, motion–environment compatibility). A motion is deemed a failure if it scores under 3 in either category or below 5 in both. As shown, HAPS 2.0 achieves higher accuracy (8.5 vs. 6.3), better compatibility (8.1 vs. 5.9), and zero failures (0 vs. 120). It also increases motion diversity (486 vs. 435) and overall annotation effort (430+ vs. 320 hours). Moreover, HAPS 2.0 refines annotation workflows and simulator design for enhanced generalization.
Altogether, HAPS 2.0 includes 26 distinct regions across 90 architectural scenes, covering 486 human activities in both indoor and outdoor contexts. Figure A2 illustrates these improvements. By offering more accurate, flexible, and diverse depictions of human actions, HAPS 2.0 provides a robust foundation for research in human motion analysis, social navigation, and beyond.
B.2 Coarse Annotation Using PSO
We adopt a coarse-to-fine strategy for positioning human motions in 3D scans. Initially, we define each region by boundary coordinates $\mathbf{B}_{\text{lo}}=(x_{\text{lo}},y_{\text{lo}},z_{\text{lo}})$, $\mathbf{B}_{\text{hi}}=(x_{\text{hi}},y_{\text{hi}},z_{\text{hi}})$, and compile an object list $\mathbf{O}=\{j_{1},j_{2},\dots,j_{n}\}$ with positions $\mathbf{p}^{j_{i}}$. We then use Particle Swarm Optimization (PSO) [25] (Algorithm A1) to locate each motion $h_{i}$ at an optimal position $\mathbf{p}^{opt}$.
Safe Distance Constraint. We set $\epsilon=1\,\mathrm{m}$ as the minimum clearance between humans and objects, ensuring a realistic layout while leaving space for agent passage.
Adaptive Penalties. Our fitness function applies penalties to placements that violate constraints (e.g., intersecting walls or overlapping humans). This strategy discourages infeasible poses and promotes plausible scene geometry alignments. The resulting coarse alignment establishes a starting point, after which we apply finer manual or semi-automated adjustments to refine multi-human interactions and ensure consistent coverage of diverse motion types.
Algorithm A1 Coarse Annotation via PSO
0:  Region $\mathbf{R}$ $\!\leftarrow\!\langle\mathbf{r},\mathbf{B}_{\text{lo}},\mathbf{B}_{\text{hi}}\rangle$, where $\mathbf{r}$ is region label and boundary coordinates $\mathbf{B}_{\text{lo}}\!=\!(x_{\text{lo}},y_{\text{lo}},z_{\text{lo}})$ and $\mathbf{B}_{\text{hi}}\!=\!(x_{\text{hi}},y_{\text{hi}},z_{\text{hi}})$; object list $\mathbf{O}\leftarrow\{j_{1},j_{2},\ldots,j_{n}\}$ with positions $\mathbf{p}_{j_{i}}\leftarrow(x_{j_{i}},y_{j_{i}},z_{j_{i}})$; human motion set $\mathbf{H}$; minimum safe distance $\epsilon\leftarrow 1~{}\text{m}$; height offset $\Delta_{z}\leftarrow 0.75~{}\text{m}$.
0:  Final positions $\mathbf{p}^{h}\leftarrow(x_{h},y_{h},z_{h})$ for each human motion $h\in\mathbf{H}$.
1:  while not all human motions placed do
2:     Filter human motions $\mathbf{H}^{\prime}\subseteq\mathbf{H}$ matching $\mathbf{r}$;
3:     Match objects $\mathbf{O}$ with human motions $\mathbf{H}^{\prime}$ based on semantic similarity to form pairs $(h_{i},j_{i})$;
4:     for each pair $(h_{i},j_{i})$ do
5:        Define search space $\mathbf{S}\leftarrow\left\langle x_{lo},x_{hi}\right\rangle\times\left\langle z_{lo},z_{hi}\right\rangle\times\left\langle y_{lo},y_{hi}\right\rangle$ around object $j_{i}$;
6:        Initialize PSO with particles randomly positioned within $\mathbf{S}$;
7:        Convergence criteria $\leftarrow$ minimal fitness change;
8:        repeat
9:           for each particle $p$ in the swarm do
10:              Compute position $\mathbf{p}^{h}$ of particle $p$;
11:              Compute fitness $f(p)$;
12:              $f(p)\leftarrow d(\mathbf{p}^{h},\mathbf{p}^{j_{i}})+P_{\text{constraints}}(p)$;
13:              where $d(\mathbf{p}^{h},\mathbf{p}^{j_{i}})$ is the Euclidean distance, and $P_{\text{constraints}}(p)$ is the penalty for constraint violations;
14:              Constraints:
15:              $d(\mathbf{p}^{h},\mathbf{p}^{j_{i}})\leq 1~{}\text{m}$; (Proximity to target object)
16:              $d(\mathbf{p}^{h},\mathbf{p}^{j_{u}})\geq\epsilon$, $\forall j_{u}\in\mathbf{O},j_{u}\neq j_{i}$; (Maintain safe distance from other objects)
17:              $\mathbf{p}^{h}\in\mathbf{R}$; (Within region boundaries)
18:              Optional: $z_{h}\geq z_{j_{i}}+\Delta_{z}$; (Height offset)
19:           end for
20:           Update particle velocities and positions using PSO update equations;
21:        until convergence criteria met
22:        Assign best particle position $\mathbf{p}^{h}$ to $h_{i}$;
23:        if no feasible solution found then
24:           Adjust PSO parameters and retry;
25:        end if
26:     end for
27:  end while
Figure A1: Multi-View Camera Setup. (a) A sample scene overview. (b) Schematic illustrating the nine camera placements around the human figure, noting key coordinates and rotations. (c) Example snapshots from the nine directional cameras, each providing a distinct viewpoint for accurate motion annotation.
Figure A2: Overview of HAPS 2.0 Scenes. These examples illustrate annotated human subjects across multiple scans in the HAPS 2.0 Dataset, highlighting a range of well-aligned motions, movements, and interactions (both with objects and with other humans).
Figure A3: Movement Examples. We present representative frames from a single set of human motions, each annotated with its corresponding movement. Activities include ascending stairs, running, and pacing. For clarity, we highlight four camera views (Cameras 2, 4, 6, 8) within the multi-camera setup to provide a comprehensive perspective of human behaviors. (Zoom in for finer details.)
B.3 Fine Annotation Using a Multi-Camera Setup
To refine the coarse placements of human motions, we draw inspiration from 3D skeleton-capture methods [22, 40] and deploy nine RGB cameras, each positioned around the motion site. As shown in Figure A1, this arrangement provides a comprehensive multi-view perspective, revealing potential collisions or misalignments between the human figure and surrounding objects.
Camera Positions & Angles. For each camera $i$ $(i=1,2,\ldots,8)$, we set its 3D location $\mathbf{p}^{\text{cam}}$ to shift by $\Delta_{x}$, $\Delta_{y}$, and $\Delta_{z}$ from the base position $\mathbf{p}^{\text{h}}$. Horizontal rotation $\theta_{\text{lr}}^{i}$ is uniformly spaced at $\frac{\pi i}{8}$, while vertical rotation $\theta_{\text{ud}}^{i}$ depends on whether $i$ is odd or even:
$\tan\theta_{\text{ud}}^{i}=\begin{cases}0,&\text{if }i\text{ is odd},\\ \frac{\Delta_{z}}{\sqrt{2}\,\epsilon},&\text{if }i\text{ is even}.\end{cases}$
For the ninth camera (overhead view), $\theta_{\text{lr}}^{9}=0$ and $\theta_{\text{ud}}^{9}=\frac{\pi}{2}$. These settings are ideal for general views and can be further adjusted in constrained spaces (e.g., narrow closets) or scenes requiring specialized viewpoints.
B.4 Fine Annotation Protocol
We adopt the following six-step procedure to fine-tune a human’s position and orientation:
  1. 1.
    Initial View. Generate an overall preview of the human figure at $\mathbf{p}^{\text{h}}$ (Figure A1(a)).
  2. 2.
    Multi-Camera Observations. Collect images from the nine cameras (Figures A1(b)–(c)). Adjust camera angles or offsets as necessary, particularly in tight scenes like small bathrooms or closets.
  3. 3.
    Vertical Collision Checks. Inspect overhead Camera 9 to detect vertical overlaps (e.g., arms interpenetrating a table). If collisions exist, identify the nearest side camera to determine how best to shift the figure.
  4. 4.
    Horizontal Translation. Modify $\Delta_{x}$ and $\Delta_{y}$ accordingly—if a nearby camera (e.g., Camera 1) reveals front-facing overlaps, shift $\mathbf{p}^{\text{h}}$ by adding or subtracting based on Camera 1’s perspective.
  5. 5.
    Side Cameras Review. Examine Cameras 2–8 to catch lingering overhang or collisions. Adjust the figure’s position proportionally, typically referencing a standard human height of 1.5 m to gauge whether shifts remain plausible.
  6. 6.
    Finalize Output. Upon confirming a collision-free layout, automatically generate the final video render and corresponding JSON metadata files.
This multi-camera process systematically eliminates misalignments, ensuring each human model remains properly integrated within the environment. The result is a more realistic portrayal of multi-human interactions and improved fidelity for downstream tasks.
"system": "You are an expert in 3D scene understanding, specializing in generating detailed multi-person interaction scenarios in complex environments. I will provide you with information about the current human motion, object positions, and the region context. Your task is to infer possible additional human interactions and describe their actions in relation to the objects and scene."
"example": ...
//examples in Listing 2
"user": Input:
[human_motion]: \"...\";
[Region]: \"...\";
[Object_1]: \"... \";
[Object_2]: \"...\";
[Object_3]: \"...\";
Based on these inputs, provide a detailed, objective description of additional people in the scene, their locations, and their interactions with objects and each other. Ensure the interactions are plausible based on the given layout and actions.
Output: ... //generation
Listing 1: Format of our LLM prompt for multi-human interactions & movements enrichment.
# Example 1
Input:
[human_motion]: "tv:An individual enjoying a sports game, cheering loudly.\"
[…]
[TV]: "\4 meters in front of the couch, placed on a white media console against the wall.\"
[…]
[table]: "\1.5 meters in front of the couch, holding some snacks and drinks.\"
Output:
"Three friends are gathered in the living room, watching a movie together. Two of them are seated comfortably on the couch, which is positioned against the wall, facing the TV. They are sharing popcorn and drinks placed on a table 1.5 meters in front of the couch. The atmosphere is relaxed, with one friend fully focused on the movie, casually eating popcorn. Another friend seated next to him. In front of the TV, a third friend stands closer, about 2 meters from the couch, is more animated, loudly cheering as they switch their attention to a sports game playing on a different screen."
Listing 2: Few-shot example for for multi-human interactions & movements enrichment.
B.5 Multi-Human Interaction & Movement Enrichment
To diversify scenes and amplify interactivity, we place additional characters into regions already featuring human motion annotations. This enables more complex interactions and varied motion trajectories. Manual insertion of extra characters, however, is time-consuming and prone to subjective bias, limiting data reliability and diversity.
Human-in-the-Loop Method. We employ large language models (LLMs) such as ChatGPT-4 and LLaMA-3-8B-Instruct to propose plausible multi-human scenarios. Each prompt integrates details about existing human motions, object positions, and regional context, guiding the LLMs to generate rich, multi-character interactions. Our prompt design uses a system prompt and few-shot examples (Listings LABEL:list:dataset_prompt and LABEL:list:few-shot-example) to ensure clarity and detail. For instance, we collect each human’s position and identify objects within 6 m, describing relative distances and orientations. The LLMs then construct additional human activities suited to the scene, merging them into cohesive multi-person narratives.
Iterative Annotation Workflow. After the LLMs produce candidate interactions, we merge outputs from ChatGPT-4 and LLaMA-3-8B-Instruct, then manually refine and validate them over four rounds [11]. This process corrects inconsistencies and ensures contextual alignment. We subsequently place new human motions according to the generated descriptions, leveraging our multi-camera technique (Sec. B.3) for precise annotation of complex activities (e.g., stair-walking, see Fig. A3).
Examples of Enriched Interactions. Figure 2 demonstrates how additional humans can populate a living room: “two people sit on the couch, sharing popcorn on a small table,” while “a third friend stands in front of the TV, cheering.” Such enriched scenes capture realistic multi-human behaviors—from casual gatherings to active cheering—offering agents a broader range of social cues for navigation and interaction.
Algorithm A2 Real-time Human Rendering in Simulation
0:  Simulation environment $\mathcal{E}$; Human motion data $\mathbf{H}$; Signal queue $\mathcal{Q}$ with maximum size $M\leftarrow 120$; Total frames $N\leftarrow 120$; Frame interval $\Delta t$.
0:  Continuous real-time rendering of $\mathbf{H}$ within $\mathcal{E}$.
1:  Initialize simulator $\mathcal{E}$, object template manager $\mathcal{T}$ in $\mathcal{E}$, human motion data $\mathbf{H}$ and signal queue $\mathcal{Q}$;
2:  Initialize total signals sent and processed to 00;
3:  // Thread 1: Signal sender thread
4:  while true do
5:     if not $\mathcal{Q}.\texttt{full}()$ then
6:        Enqueue signal “REFRESH_HUMAN” into $\mathcal{Q}$;
7:        Increment total signals sent;
8:     end if
9:     Sleep for $\Delta t$;
10:  end while
11:  // Thread 2: Main thread
12:  while simulation is running do
13:     if new episode starts then
14:        Clear $\mathcal{Q}$ and reset total signals sent to 00;
15:        Remove previous human models from $\mathcal{E}$;
16:     end if
17:     // Agent handles signals before observation
18:     while not $\mathcal{Q}.\texttt{empty}()$ do
19:        Dequeue signal from $\mathcal{Q}$;
20:        $t\leftarrow(\text{total signals processed})\bmod N$Compute current frame index};
21:        Remove previous human models from $\mathcal{E}$;
22:        for each human motion $h\in\mathbf{H}$ do
23:           Retrieve motion category, translation, and rotation of $h$ at frame $t$;
24:           Load template $\tau_{h}$ into $\mathcal{T}$;
25:           Add human $o_{h}$ to $\mathcal{E}$ using template $\tau_{h}$;
26:           Set translation and rotation of $o_{h}$;
27:        end for
28:        Increment total signals processed;
29:     end while
30:     Agent observes environment and makes decision;
31:  end while
B.6 Real-Time Human Rendering
We integrate dynamic human models into the simulation through a multi-threaded pipeline inspired by Producer–Consumer principles and Java-style signaling (check details in Algorithm A2). This setup enables agents to observe and respond to human motions in real time, facilitating adaptable navigation policies.
System Initialization. We begin by loading the environment $\mathcal{E}$, the set of human motions $\mathbf{H}$, and an object template manager $\mathcal{T}$ that handles 3D model templates efficiently.
Signal Sender Thread (Thread 1). At intervals $\Delta t$, Thread 1 places “refresh” signals into a queue $\mathcal{Q}$. If $\mathcal{Q}$ is full, it pauses until earlier signals are processed, preventing data overload. This thread models a continuous stream of human motion updates at a fixed frequency.
[…]
Synchronization & Consistency. We refresh human models immediately before the agent’s perception step, ensuring it observes the latest motion state. Upon starting a new episode, $\mathcal{Q}$ is cleared, and signal counters reset, so human motions revert to frame 0, maintaining consistency across episodes. This real-time process keeps human activities synchronized with the agent’s action cycle, creating dynamically evolving scenes where agents must adapt to changing bystander locations and behaviors.
Figure A4: The visualization of Human Counting.
Figure A5: Statistics on human motion annotations in HA-VLN simulator: (a) Distribution of humans by region; (b) Top 15 motions without multi-human interaction & movement enrichment; (c) Top 15 motions with enrichment; (d) Distribution of human trajectory lengths (in meters); (e) Impact of human presence on the environment, categorized as direct, indirect, and no impact. (Zoom in to view)
B.7 API Design
Discrete Environment (DE). In our discrete setting, all agent and human positions are tracked via a real-time navigational graph displayed in a 2D top-down view. Each human’s activity is stored as a tuple $\langle p_{h},d_{agent},\theta_{relative},a_{status}\rangle$, where $p_{h}$ is the human’s 2D coordinate, $d_{agent}$ is the distance to the agent, $\theta_{relative}$ is the relative orientation, and $a_{status}$ indicates the activity state. This representation supports efficient, simultaneous tracking of multiple humans in a discrete viewpoint space.
Multi-Entity Detection & Tracking. We employ object detection on each discrete panorama to identify humans, assigning unique IDs for continuous monitoring throughout the navigation process. By linking recognized human poses to specific graph nodes, we anchor their activities to well-defined spatial references.
User Interface. A specialized UI presents a bird’s-eye view of the 2D graph, allowing researchers to visualize, annotate, and adjust human behaviors in real time. This interface significantly streamlines data annotation and analysis for discrete human-aware navigation research.
Continuous Environment (CE). Our API in continuous mode mainly focuses on three components: (1) Human Activity Monitoring, (2) Environmental Perception, and (3) Navigation Support.
(1) Human Activity Monitoring. We track and analyze human activity in real time as in Sec. 3. When collisions occur, the agent reverts to its prior position, and we identify whether the obstacle is human or an inanimate object. For human collisions, we log the coordinates and motion state to inform potential reward-shaping strategies. Distance and orientation estimates derive from agent–human coordinate data. For instance, we employ the Grounding-DINO [34] detector on RGB inputs with the prompt “human” to count individuals. Figure A4 illustrates how human detection bounding boxes enable real-time counting.
(2) Environmental Perception. We maintain a dynamic scene graph comprising static elements (e.g., buildings, furniture) and moving humans. The agent continuously updates this graph by fusing positional changes and human motion data in its vicinity. This ensures real-time awareness of human activities for downstream decisions.
(3) Navigation Support. An A*-based planner computes candidate trajectories while accounting for both dynamic humans and static obstacles. During execution, we monitor any divergence between the agent’s chosen route and the planner’s recommended path. This method highlights human-centric obstacles and informs the agent’s short-term re-planning steps. Our unified API supports real-time detection, tracking, and socially compliant navigation decisions in both discrete and continuous modes. It simplifies multi-human scene management, ensures intuitive collision handling, and provides robust path-planning assistance—together forming a foundation for advanced human-aware navigation algorithms.
B.8 Human Activities Annotation Data Analysis
Human Distribution by Region. Figure A5(a) illustrates how 910 human figures are dispersed across 90 buildings, averaging about nine individuals per building. Even though each person moves independently, this distribution ensures robust and dynamic multi-human interactions, closely mirroring real-world scenarios.
Motion Frequency Analysis. Figures A5(b)–(c) compare the 15 most frequent motions before and after multi-human enrichments. While the total number of motions increases, we also embed additional movement patterns and group interactions into existing actions. For instance, “talking on the phone while pacing” may now involve extended pacing distances or layered scenarios like “a couple having a quiet conversation” or “friends sharing stories over dinner.”
Movement Distance Analysis. Figure A5(d) displays the distribution of trajectory lengths for actively moving humans. Specifically, 22.4% cover distances up to 0.5 m, 37.3% reach 0.5–1 m, 25.0% span 1–1.5 m, 11.6% extend 1.5–2 m, and the remaining 3.7% exceed 2 m. This wide range reflects the diverse indoor and outdoor behaviors encompassed in the dataset.
Human Impact Analysis. As shown in Figure A5(e), humans exert a notable influence on navigation paths: 35.5% of the 16,844 paths in HA-VLN physically intersect with human motion, while 49.7% of viewpoints are indirectly affected (i.e., humans are visible along the route). These statistics underline the importance of accounting for human presence and movement trajectories when designing real-world navigation agents.
Table A4: Instruction Samples from the HA-R2R Dataset. Text in purple highlights human-related actions/movements, while text in blue indicates explicit agent-human interaction cues. These examples illustrate how HA-R2R integrates dynamic human considerations and social awareness into navigation instructions.
1. Exit the library and turn left. As you proceed straight ahead, you will enter the bedroom, where you can observe a person actively searching for a lost item, perhaps checking under the bed or inside drawers. Continue moving forward, ensuring you do not disturb his search. As you pass by, you might see a family engaged in a casual conversation on the porch or terrace, be careful not to bump into them. Maintain your course until you reach the closet. Stop just outside the closet and await further instructions.
2. Begin your path on the left side of the dining room, where a group of friends is gathered around a table, enjoying dinner and exchanging stories with laughter. As you move across this area, be cautious not to disturb their gathering. The dining room features a large table and chairs. Proceed through the doorway that leads out of the dining room. Upon entering the hallway, continue straight and then make a left turn. As you walk down this corridor, you might notice framed pictures along the walls. The sound of laughter and conversation from the dining room may still be audible as you move further away. Continue down the hallway until you reach the entrance of the office. Here, you will observe a person engaged in taking photographs, likely focusing on capturing the view from a window or an interesting aspect of the room. Stop at this point, ensuring you are positioned at the entrance without obstructing the photographer’s activity.
3. Starting in the living room, you can observe an individual practicing dance moves, possibly trying out new steps. As you proceed straight ahead, you will pass by couches where a couple is engaged in a quiet, intimate conversation, speaking softly to maintain their privacy. Continue moving forward, ensuring you navigate around any furniture or obstacles in your path. As you transition into the hallway, notice another couple enjoying a date night at the bar, perhaps sharing drinks and laughter. Maintain a steady course without disturbing them, keeping to the right side of the hallway. Upon reaching the end of your path, you will find yourself back in the living room. Here, a person is checking their appearance in a hallway mirror, possibly adjusting their attire or hair. Stop by the right candle mounted on the wall, ensuring you are positioned without blocking any pathways.
4. Begin by leaving the room and turning to your right. Proceed down the hallway, be careful of any human activity or objects along the way. As you continue, look for the first doorway on your right. Enter through this doorway and advance towards the shelves. Once you reach the vicinity of the shelves, come to a halt and wait there. During this movement, avoid any obstacles or disruptions in the environment.
Appendix C Agent Details
C.1 HA-R2R Instruction Examples
Table A4 illustrates four sample instructions from the Human-Aware Room-to-Room (HA-R2R) dataset. These examples encompass multiple scenarios: multi-human interactions (e.g., 1, 2, 3), direct agent–human encounters (e.g., 1, 2, 3), situations with four or more bystanders (e.g., 3), and paths devoid of humans (e.g., 4). Together, they demonstrate how HA-R2R challenges an agent with diverse human-aligned instructions.
C.2 HA-R2R Instruction Generation
To create enriched instructions for HA-R2R, we use ChatGPT-4o and LLaMA-3-8B-Instruct to expand upon R2R-CE’s original textual data. Our strategy involves a carefully crafted few-shot prompt, combining a system prompt (Listing LABEL:list3:enrichment_prompt) and few-shot examples (Listing LABEL:list4:enrichment-few-shot-example).
Prompt Structure. The system prompt lays out guidelines for generating instructions that emphasize social context. It encourages mentioning human activities and interactions relevant to navigation paths. Few-shot examples then illustrate the desired format, including references to human behavior (e.g., “someone quietly making a phone call; keep your voice down as you proceed”), positional references, and object interactions.
Iterative Refinement. In early trials, the models sometimes produced extraneous or subjective content, lacking sufficient detail on human activities. We iteratively refined the system prompt and examples, clarifying the need for neutral tone, accuracy, and contextual alignment with human-related scenarios. In each round, we analyzed model outputs, identified discrepancies, and adjusted examples to showcase more detailed, coherent, and socially aware instructions. This process guided ChatGPT-4o and LLaMA-3-8B-Instruct toward generating instructions that fully integrate human-centric elements—such as bystander activities, relevant spatial cues, and subtle behavioral recommendations. The final HA-R2R instructions thus reflect enriched scene descriptions where agents must account for diverse, real-world nuances involving human presence.
"system": "You are tasked with generating an objective and detailed description of a path and the related scene information based on the given input. Focus solely on observable entities and actions in the scene, avoiding descriptions of smell, atmosphere, or subjective feelings. Ensure clarity and precision in object and human activity descriptions."
"examples": ... //examples in Listing 4
"user": Input:
[instruction_text]: \"...\";
[start]: \"...\";
[during]: \"...\";
[end]: \"...\";
Please provide an objective, step-by-step description of the path, focusing on the objects and human activities observed at each stage of the navigation (beginning, during, and end). Expand your description by including character actions, objective descriptions of objects, and details of the movement process. Your description should maintain a neutral tone and avoid embellishments. Please simulate the robots movement along the path, while incorporating possible connections or interactions between the robot, objects, and human activities.
Output: ... //generation
Listing 3: Format of our LLM prompt for HA-R2R instrcution generation.
# Example 1:
Input:
[instruction_text]: \"Go through the doorway to the hall. Go down the hall past the landing. Take the first door on the right and stop in the doorway. \";
[start]: \"other room:A person using the room for a quiet phone call. \";
[during]: \"other room:A person using the room for a quiet phone call. \";
[end]: \"hallway:An individual cleaning or vacuuming. \";
Output:
You should start from the room. Before you walk through the doorway into the hallway, you will notice someone quietly making a phone call, so please remain quiet as you move. Then, proceed down the hallway past the landing. As you continue through the corridor, you will see a responsible cleaner using a vacuum to tidy up. Finally, your destination is at the end of the hallway, enter the first door on the right and stop in the doorway.
Listing 4: Few-shot examples for HA-R2R instrcution generation.
C.3 HA-R2R Data Analysis
Word Frequency Analysis. We conduct a word frequency study on HA-R2R to gauge its capacity for representing realistic, human-centric scenarios. Figures A6(a) and (b) illustrate frequently used nouns and verbs, respectively, confirming the dataset’s focus on both spatial navigation and social interactions.
Nouns. The five most common nouns are room, hallway, turn, area, and path, with room alone appearing over 15,000 times. Other notable terms (person, doorway, kitchen) highlight spatial complexity and social elements such as conversation, activities, and someone.
Verbs. The five most frequent verbs—is, continue, proceed, ensuring, be—reveal an action-oriented narrative, while additional terms (engaged, observe, notice, avoid, maintain) underscore instructions geared toward social awareness and precise route-following.
Human Impact Analysis. Figure A6(c) shows that most instructions contain 20–60% human-related content, reflecting the dataset’s emphasis on people in everyday scenes.
Comparisons of word clouds in Figures A6(d) and (e) confirm that while both human-aligned and non-human segments use common navigational verbs (walk, left, right), instructions involving humans introduce additional social context (couple, man, painting). This integration of interpersonal cues elevates HA-R2R beyond simple route directives, better mirroring real-world navigation challenges in human-filled environments.
Figure A6: Statistics for the HA-R2R Dataset. (a) Verb frequency distribution for all instructions. (b) Noun frequency distribution for all instructions. (c) Distribution of human impact within HA-R2R (originally Fig. A4 in the main text; figure numbering differs due to inserted figures). (d) Word cloud of instructions not aligned with human activities. (e) Word cloud of instructions explicitly involving human actions. Larger font size indicates higher frequency or proportion in the dataset.
C.4 Visual and Depth Embeddings
Following VLN-CE [28], we employ parallel streams to process RGB and depth images. Each viewpoint produces a set of features from two specialized ResNet-50 models:
  • RGB Features. Let $\{v^{rgb}_{1},v^{rgb}_{2},\dots,v^{rgb}_{k}\}$, where $v^{rgb}_{i}\in\mathbb{R}^{2048}$, be outputs of a ResNet-50 pretrained on ImageNet.
  • Depth Features. Let $\{v^{d}_{1},v^{d}_{2},\dots,v^{d}_{k}\}$, where $v^{d}_{i}\in\mathbb{R}^{128}$, be outputs of another ResNet-50 pretrained on Gibson-4+ [53] and MP3D for point-goal navigation.
We fuse these two feature streams along with a directional encoding $d_{i}$ indicating spatial orientation:
$v_{i}=\bigl{[}\;v^{rgb}_{i}\,W_{\mathrm{rgb}};\;\,v^{d}_{i}\,W_{\mathrm{depth}};\;\,d_{i}\;\bigr{]}\,W_{\mathrm{merge}},$
where $W_{\mathrm{rgb}}$, $W_{\mathrm{depth}}$, and $W_{\mathrm{merge}}$ are learnable projection matrices with ReLU activation. The directional encoding $d_{i}$ is constructed by repeating $(\cos\theta^{i}_{t},\;\sin\theta^{i}_{t})$ 32 times, where $\theta^{i}_{t}$ measures the relative heading offset of the agent. The fused embedding $v_{i}\in\mathbb{R}^{d}$ is either 512 or 768 dimensions, matching the requirements of our HA-VLN-CMA or HA-VLN-VL agent, respectively. Both ResNet backbones remain fixed during training, ensuring consistent and stable representations from the RGB and depth channels throughout the learning process.
C.5 Text Embeddings
For the HA-VLN-VL agent, we utilize text embeddings from PREVALENT [18], which was pre-trained on 6.58M image–text–action triplets, thereby capturing broad contextual cues for navigation. Conversely, the HA-VLN-CMA agent adopts embeddings from BERT [10], also widely used for its strong language representations.
Formally, let =$\ell=\{w_{1},\dots,w_{n}\}$ be a sequence of tokens representing the instruction. Each token $w_{i}$ is mapped to a one-hot vector $e_{i}\in\mathbb{R}^{V}$, where $V$ is the vocabulary size. An embedding matrix $E\in\mathbb{R}^{V\times d}$ then projects $e_{i}$ into a continuous $d$-dimensional space:
$x_{i}=E^{\top}\,e_{i},\quad x_{i}\in\mathbb{R}^{d}.$
In this manner, each discrete token $w_{i}$ is transformed into a trainable embedding $x_{i}$, forming the foundation of the model’s linguistic understanding.
Figure A7: Network Structures. (a) HA-VLN-VL adopts a BERT-like transformer with a specialized state token. RGB and depth inputs are compressed by ResNet-50 and concatenated, while instruction tokens feed a BERT-like encoder. A multi-layer transformer computes cross-modal attention, producing per-step action probabilities via average-pooling and a final projection. In both architectures, continuous or discrete commands are then derived for navigation based on the agent’s policy output. (b) HA-VLN-CMA employs a cross-modal attention (CMA) module combined with a GRU policy. RGB and depth images are first processed by two ResNet-50 encoders and fused into a single feature stream, which attends to the instruction tokens; the fused features are then fed into a GRU and MLP to predict actions.
Figure A8: Trajectory Comparison Under Human vs. No-Human Conditions. We illustrate the same episode’s trajectories predicted by BEVBert [1], trained on VLN-CE, in scenarios with (bottom) and without (top) human presence. In the top row, no bystanders are present, and the agent follows its instructions with minimal collision risk. In the bottom row, bystanders and human-aligned cues lead to altered motion decisions, sometimes creating additional collision challenges or deviations.
Figure A9: Agent Trajectory Examples (HA-VLN-CMA). The top row depicts a failed navigation case in which the agent fails to avoid an oncoming human, resulting in a collision. By contrast, the bottom row illustrates a successful run: the agent proactively adjusts its path to the right, thereby avoiding interference and completing the task collision-free.
C.6 HA-VLN-VL Structure
Model Overview. HA-VLN-VL adopts a BERT-like architecture inspired by Recurrent VLN-BERT [20], extending it to handle human-aware navigation. At each timestep $t$, the model receives the previous state $s_{t-1}$, language tokens $X$, and fused RGB–depth visual features $V_{t}$ (Sec. C.4). It outputs an updated state $s_{t}$ and an action distribution $p_{t}^{a}$:
$s_{t},\;p_{t}^{a}=\mathrm{HA\text{-}VLN\text{-}VL}\bigl{(}s_{t-1},\,X,\,V_{t}\bigr{)}.$
State Token. In line with BERT conventions, the model maintains a state token $s_{t}$ that encapsulates the agent’s internal context. Initially, $s_{0}$ is set to the embedding of the [CLS] token. At each step, the state token is updated by appending the agent’s previously executed action $a_{t}$ and projecting the resulting vector:
$s_{t}=[\,s^{\prime}_{t};\,a_{t}\,]\;W_{s},$
where $s^{\prime}_{t}$ is the final Transformer-layer output, and $W_{s}$ is a learnable projection matrix.
Visual Attention. To decide the next action, we compute attention scores between $s_{t}$ and the set of visual tokens $V_{t}=\{v_{1},v_{2},\dots,v_{n}\}$, covering navigable directions plus a “stop” option:
$A_{s,v}^{t}=\mathrm{Softmax}\!\Bigl{(}\frac{Q_{s}\,K_{v}^{\top}}{\sqrt{d_{h}}}\Bigr{)},$
where $Q_{s}$ is derived from $s_{t}$ and $K_{v}$ from $v_{i}\in V_{t}$. The model then aggregates these attention scores via an average-pooling step:
$p_{t}^{a}=\mathrm{\overline{AveragePool}}\bigl{(}A_{s,v^{\prime}}^{t}\bigr{)},$
yielding an action distribution over possible moves. The agent selects:
$a_{t}=\arg\max\bigl{(}p_{t}^{a}\bigr{)}.$
Training Objective. HA-VLN-VL is optimized through a combination of supervised imitation learning—to mimic ground-truth trajectories—and optional reinforcement learning, which rewards safe and efficient paths. As depicted in Figure A7(a), the model continuously refines its understanding of language instructions and visual cues, offering robust and socially aware navigation.
C.7 HA-VLN-CMA Structure
Architecture Overview. HA-VLN-CMA is a dual-stream visual-language agent featuring Cross-Modal Attention (CMA) and a recurrent decoder for navigation in human-populated scenarios (see Figure A7(b)). It processes two visual channels—RGB and Depth—alongside language instructions, then outputs an action at each time step.
Dual-Stream Visual Encoding. Following Sec. C.4, each observation $o_{t}$ is split into:
$v^{\mathrm{rgb}}_{t}=\mathrm{ResNet}^{\mathrm{rgb}}(o_{t}),\quad v^{\mathrm{d}}_{t}=\mathrm{ResNet}^{\mathrm{depth}}(o_{t}),$
where $\mathrm{ResNet}^{\mathrm{rgb}}$ and $\mathrm{ResNet}^{\mathrm{depth}}$ are separate backbones for RGB and Depth, respectively. The fused feature representation is
$v_{i}=\bigl{[}v^{\mathrm{rgb}}_{i}W_{\mathrm{rgb}};\;v^{d}_{i}W_{\mathrm{depth}};\;d_{i}\bigr{]}\,W_{\mathrm{merge}},$
where $W_{\mathrm{rgb}}$, $W_{\mathrm{depth}}$, and $W_{\mathrm{merge}}$ are projection matrices, and $d_{i}$ is a direction encoding (Sec. C.4).
Language Encoder. Textual instructions $\{w_{1},\dots,w_{T}\}$ are transformed into contextual embeddings
$l=\mathrm{BERT}(w_{1},\dots,w_{T}).$
These embeddings capture the semantic structure of the instruction and serve as input to the cross-modal module.
Cross-Modal Attention & Recurrent Decoding. At time step $t$, we attend to the language features using multi-head attention:
$\hat{f}_{t}=\mathrm{MultiHeadAttn}(v_{t},\;l),$
where $\mathrm{Attention}(Q,K,V)=\mathrm{softmax}\!\bigl{(}\frac{QK^{\top}}{\sqrt{d_{k}}}\bigr{)}V.$ Multi-head attention helps handle lengthy and detailed instructions by learning multiple representations in parallel.
Next, we combine the resulting multimodal embeddings with the previous action $a_{t-1}$ in a GRU-based decoder:
$f_{t}=\mathrm{GRU}\bigl{(}\,[\,(\,v_{t},l),\,a_{t-1}\,],\,f_{t-1}\bigr{)},$
where $f_{t-1}$ is the previous hidden state. Finally, an MLP outputs the action distribution:
$a_{t}=\mathrm{softmax}(\mathrm{MLP}(f_{t})),$
where $\mathrm{MLP}(f_{t})=W_{a}\,f_{t}+b_{a}$, and $a_{t}$ is sampled from $P(a_{t}|f_{t})$.
Training Objectives. HA-VLN-CMA is trained end-to-end with a mixture of imitation learning (to mimic ground-truth paths) and reinforcement learning (to encourage collision-free, socially compliant navigation). By learning from both paradigms, the agent refines its ability to balance path efficiency and safe distancing in human-populated environments.
Appendix D Experiments Details
D.1 Evaluation Metrics
We adopt a two-tier evaluation protocol for HA-VLN, measuring both perception (human awareness) and navigation (task completion). Perception metrics track how effectively the agent detects and responds to dynamic humans, while navigation metrics assess overall performance.
Total Collision Rate (TCR). Given the strong impact of human activities around critical nodes (viewpoints), we manage dynamic humans to ensure precise measurement. For navigation instance $i$, let $A^{c}_{i}$ be the set of human activities at these critical nodes. We define:
$\mathrm{TCR}=\frac{\sum_{i=1}^{L}(\,c_{i}-|\,A^{c}_{i}|\,)}{L},$
where $c_{i}$ counts collisions within 1 m of a human. TCR quantifies how often collisions occur in human-occupied zones.
Collision Rate (CR). CR is the fraction of navigation instances incurring at least one collision, conditioned on the fraction $\beta$ of instructions influenced by humans:
[…]
Unlike TCR, CR highlights whether a collision occurred at all—offering insight into safety over entire trajectories.
Navigation Error (NE). NE is the mean distance between the agent’s final position and the intended target:
$\mathrm{NE}=\frac{\sum_{i=1}^{L}d_{i}}{L},$
where $d_{i}$ is the agent–target distance at episode end.
Success Rate (SR). SR measures the ratio of episodes completed with zero collisions:
$\mathrm{SR}=\frac{\sum_{i=1}^{L}\mathbb{I}\bigl{(}c_{i}-|A^{c}_{i}|=0\bigr{)}}{L},$
where $\mathbb{I}$ is 1 if the agent avoids collisions, and 0 otherwise.
D.2 Ground Truth Path Annotation
In HA-VLN-CE, the agent must reach within 3 m of the target while minimizing collisions. To label ground-truth paths, we use an A*-based heuristic search that identifies the shortest viable route, dynamically re-planning when obstacles block progress.
D.3 Visualization of Navigation
Figures A8A9 illustrate trajectories predicted by BEVBert [1] (trained on VLN-CE) and HA-VLN-CMA, which showcases success and failure in human-filled or empty environments.
Failures with Human Crossing. In Figure A8, the agent performs well when no bystanders are present. Yet in a human-populated setting, it fails to adjust at step 37 when a volunteer crosses its path, leading to collision.
Collision vs. Avoidance. Figure A9 similarly shows two scenarios. At step 39 in the top pane, a direct approach by a bystander overwhelms the agent, causing a collision. In the bottom pane at step 22, the agent successfully deviates upon sensing a person nearby, avoiding any collision altogether. These visualizations confirm that dynamic human presence greatly complicates navigation, highlighting the need for robust social-aware models.
Figure A10: Unitree GO2-EDU quadruped robot used in real-world experiments. The platform integrates an NVIDIA Jetson Orin NX for onboard computation, an IMU for pose estimation, an Intel RealSense D435i camera for visual perception, and a 3D LiDAR for obstacle/human detection.
Figure A11: Navigation success in an office setting (left: no humans, right: with humans). Top: The given instruction for the robot. Middle: A third-person view of the robot’s path. Bottom: The robot’s egocentric camera feed.
Figure A12: Navigation failure in an office setting. A volunteer abruptly changes position, causing the robot to collide mid-path. This scenario highlights the difficulty of adapting to sudden human movement in confined workspaces.
Figure A13: Navigation success in a living room (left: no bystanders, right: with bystanders). The robot follows instructions toward the sofa and dining area, keeping safe distances while navigating around volunteers.
Figure A14: Navigation failure in a living room with multiple bystanders. Attempting to move beyond the sofa toward a dining area, the robot collides with a volunteer who abruptly stands and shifts position. This underscores how unpredictable human motion can disrupt the agent’s intended path, requiring rapid re-planning.
Figure A15: Navigation success in a hallway (left: no bystanders, right: with bystanders). When volunteers appear, the robot halts or deviates to avoid collisions, showcasing adaptive behavior in a constrained corridor.
Figure A16: Navigation failure in a hallway. A volunteer’s sudden positional change causes a mid-path collision and mission failure, reflecting the challenge of unpredictable human movement, even in comparatively open corridors.
D.4 Validation on Real-World Robots
To deploy our navigation agents on physical hardware, the robot is equipped with an NVIDIA Jetson NX for AI inference and a Raspberry Pi 4B for motion control. The Jetson handles core navigation computations (receiving camera images and inferring action commands), while the Pi executes high-level movement directives such as turn left or move forward. We set a minimum step size of 0.25 m and a rotation increment of 15 degrees. An onboard IMU continuously monitors the robot’s orientation and position, ensuring movement commands align with issued directives.
Robot Setup. Our evaluations use a Unitree GO2-EDU quadruped (Figure A10), featuring an Intel Realsense D435i camera providing RGB imagery at a $90^{\circ}$ field of view and a 3D LiDAR below the camera for obstacle detection. The IMU refines positional and orientational control, enabling consistent motion execution.
Visual Demonstrations. Figures A11, A13, and A15 show the robot traversing distinct indoor environments—offices, living rooms, and hallways—guided by natural-language instructions. In Figure 5, the robot navigates around multiple people, leveraging camera inputs to avoid collisions through minor path adjustments. Although the agent typically succeeds in reaching its destination, collisions remain possible when bystanders change their trajectories unexpectedly. Figures A12, A14, and A16 illustrate such scenarios, highlighting real-time challenges in unpredictable, human-inhabited spaces. Additional demonstrations, including a compilation video on our project webpage, further illustrate the robot’s performance and underscore how human-aware training significantly aids sim-to-real transfer in dynamic indoor environments.